# Happy family of stable marriages

### 09 Apr 2026 v2

@@ 1 Introduction
where $\nu_{x}$ is the unit normal to $\partial K$ at $x$,
@@
is the volume of the unit (n1)(n-1)-ball in Euclidean space. Considering a centrally symmetric convex set $B\in\mathbb{R}^{n}$, we are looking for a sensible definition of a perimeter of a set $K\subset\mathbb{R}^{n}$ corresponding to the norm $\|\cdot\|_{B}$ induced by $B$ as a unit ball. Our definition is inspired by the affine invariant definition of Busemann (1) in dimension 22. In fact, if we consider $K$ to be the unit ball $B$ in $\mathbb{R}^{2}$, then equation (1) is the actual perimeter of $B$ in the metric $\|\cdot\|_{B}$ (see Sec. 3), called the self-perimeter of $B$.
In our approach, we replace $\omega^{(n-1)}$ in $\sigma_{K}(\nu_{x})$ (2) by a recursive definition of the volume $\omega^{(n-1)}_{B}(\nu)$, depending on $B$ and on the unit vector $\nu\in\mathbb{R}^{n}$. See Definition 4.1 below.
@@
$|\partial K|_{B}:=\int_{\partial K}\omega^{(n-1)}_{B}(\nu_{x})d{\cal H}_{n-1}(x)\$
while its affine invariant perimeter is an adaptation of (1, 2):
$P(K)=\int_{\partial K}\frac{\omega^{(n-1)}_{B}(\nu_{x})}{{\cal H}_{n-1}(K\cap\Sigma^{n-1}(\nu_{x}))}d{\cal H}_{n-1}(x)\ ,$
This definition coincides with Busemann’s definition for $n=2$ since $\omega^{(1)}=2$ is independent of direction in $\mathbb{R}^{2}$. However, it replaces Busemann’s preservation of the Euclidean volume of the unit ball with the preservation of the Euclidean ratio $P(B)/\omega_{B}^{(n)}=n$ of perimeter/volume of the unit ball in $n$ dimensions (Definition 4.1).
In 2D, the uniquely defined perimeter (measured via the norm) sets the area base case as $V_{2}=P_{2}/2$. For dimensions $n\geq 3$, the surface area is defined by integrating over the Euclidean sphere $S^{n-1}$, normalizing each surface element by the (n1)(n-1)-dimensional volume of the unit ball’s central cross-section orthogonal to the specified direction. It is remarkable that while Busemann’s definition (1) does not show any connection with the Minkowski norm, it does give the self-perimeter when applied to the unit ball of the same norm in dimension 2. This seems to be just a coincidence, due to the ansatz $V(B^{(1)})=\omega^{(1)}=2$.
As we will show, the exact cancellation of the Jacobian in this formulation ensures that this definition of volume is strictly preserved under origin-preserving affine transformations. We extend the definition of $\omega_{B}^{(n)}$ to non-centrally symmetric convex bodies $B$, calculate explicitly the self-volume/perimeter of an $n$-simplex $\Delta^{n}$, and prove its convexity with respect to the chosen center. This implies the convexity of the self-perimeter/volume of any convex body in $\mathbb{R}^{n}$ composed of Cartesian products of simplices, 1D segments, and 2D convex sets. At the end of Section 4 we pose the following conjecture (see Section 4.4):
Conjecture
@@ 3 Self-perimeter in the 2-plane
where dsds is the Euclidean length of the boundary of $B$, and $\vec{\tau}(s)$ is the tangent vector to the boundary at $s\in\partial B$. An explicit expression in terms of an integral on the Euclidean circle $\mathbb{S}^{1}$ is:
@@
Hence, $r_{K,p}(v)$ is a concave function of pp.
@@
Since the self-perimeter $P(p)$ is an integral (a continuous sum) of strictly convex functions over the boundary $\partial K$ with respect to a positive length measure dsds, the function $P(p)$ itself is strictly convex on $\text{int}(K)$.
@@ 4.1 Examples
In this example, we study the self-volume of a non-centrally symmetric ball. Let $\Delta^{n}$ be an $n$-dimensional simplex with $n+1$ vertices $v_{1},\ldots,v_{n+1}\in\mathbb{R}^{n}$. Any interior point $p\in\text{int}(\Delta^{n})$ can be uniquely represented by its strictly positive barycentric coordinates $\lambda=(\lambda_{1},\dots,\lambda_{n+1})$, where $p=\sum_{i=1}^{n+1}\lambda_{i}v_{i}$ and $\sum_{i=1}^{n+1}\lambda_{i}=1$, with $\lambda_{i}>0$ for all ii. Let $F_{i}$ denote the (n1)(n-1)-dimensional face opposite to the $i$-th vertex.
@@
where $\{k_{1},\ldots,k_{n-1}\}\subset\{1,\ldots,n+1\}$ are n1n-1 different indices ($k_{i}\neq k_{j}$ for $i\neq j$) and $\lambda_{1},\ldots,\lambda_{n+1}$ are the barycentric coordinates of pp.
@@ 4.2 Convexity of self-volume
The self-volume $\omega^{(n)}(\Delta^{n}-p)$ is a strictly convex function of the center pp. In particular, there exists a unique $p^{*}\in\Delta^{n}$ for which $\omega^{(n)}(\Delta^{n}-p^{*})$ attains its minimum, and $\lambda_{i}=\frac{1}{n+1}$, $i=1,\ldots,n+1$ are the barycentric coordinates of $p^{*}$.
@@
Let $n=1$ and $m=1$. Both AA and $B$ are one-dimensional segments, denoted $I_{1}$ and $I_{2}$. The Cartesian product $I_{1}\times I_{2}$ is a 2D rectangle. The boundary $\partial(I_{1}\times I_{2})$ decomposes into two parts: $(\partial I_{1}\times I_{2})\cup(I_{1}\times\partial I_{2})$. For the faces $\partial I_{1}\times I_{2}$, the normal vectors are parallel to $I_{1}$. The orthogonal cross-section is $I_{2}$. The Euclidean length of these two faces is $2|I_{2}|$. The contribution to the self-perimeter is:
@@
By the recursive definition, the remaining integral is exactly the self-perimeter of AA, denoted $P^{(n)}(A)$. Since $P^{(n)}(A)=n\omega^{(n)}_{A}$, we obtain:
@@ 5.1 Background and Motivation
In classical convex geometry, the Minkowski and Alexandrov problems seek to reconstruct a convex body from its standard Euclidean surface area measure or integral Gauss curvature, respectively. In this framework, we introduce a novel variation: an Alexandrov-type problem formulated with respect to the self-surface measure of a centrally symmetric convex set. Instead of relying on the Euclidean surface area, the perimeter and volume are defined intrinsically. For a centrally symmetric convex set $K\subset\mathbb{R}^{n}$, the self-perimeter is constructed recursively by integrating over the essential boundary, normalized by the self-volume of the (n1)(n-1)-dimensional cross-sections. The inverse problem therefore asks: Given a measure $\mu$ on $S^{n-1}$, does there exist a centrally symmetric convex set $B\subset\mathbb{R}^{n}$ whose self-surface measure is exactly $\lambda\mu$?
@@
$\int_{\{s\in\mathbb{S}^{1},\phi^{(p)}(s)+\phi_{0}>0\}}\sqrt{\phi^{(p)}(s)+\phi_{0}}\,ds-\int_{\{s\in\mathbb{S}^{1},\phi^{(p)}(s)+\phi_{0}<0\}}\sqrt{-(\phi^{(p)}(s)+\phi_{0})}\,ds=0.$ (20)
@@ 6 Conclusion
In this paper, we introduced a natural, recursive definition for volume and surface area in $n$-dimensional normed spaces, centered on preserving the Euclidean ratio $P(B)/\omega^{(n)}_{B}(B)=n$ [2, 3]. We demonstrated that this construction is invariant under origin-preserving affine transformations and polar duality, satisfying fundamental geometric requirements for measures in Minkowski spaces [1, 5]. For the two-dimensional case, we utilized the property that the self-perimeter is a strictly convex functional with respect to the interior reference point, which ensures the existence of a unique minimizing generalized centroid for any convex body [8]. We further discussed the established result that the minimum self-perimeter of any planar convex body is bounded above by 9, with the maximum value attained if and only if the body is a triangle evaluated at its centroid [7]. This sharp upper bound complements the classical results for centrally symmetric sets [6]. Finally, we explored an Alexandrov-type problem for the associated self-surface measure [4]. Through a perturbative analysis around the Euclidean disc, we identified a rigidity result for the surface measure $\mu$, revealing a significant obstruction to solvability: a necessary condition for the existence of a solution is the absence of certain harmonic components in the density perturbation. These findings open new avenues for research, particularly regarding the global solvability of the inverse problem and the extremal values of the self-volume in higher dimensions.

### 13 Apr 2026 v3

@@ 1 Introduction
where $\nu_{x}$ is the unit normal to $\partial K$ at $x$.
@@
is the volume of the unit (n1)(n-1)-ball in Euclidean space. Considering a centrally symmetric convex set $B\in\mathbb{R}^{n}$, we are looking for a sensible definition of a perimeter of a set $K\subset\mathbb{R}^{n}$ corresponding to the norm $\|\cdot\|_{B}$ induced by $B$ as a unit ball. Our definition is inspired by the affine invariant definition of Busemann (1) in dimension $2$. In fact, if we consider $K$ to be the unit ball $B$ in $\mathbb{R}^{2}$, then equation (1) is the actual perimeter of $B$ in the metric $\|\cdot\|_{B}$ (see Sec. 3), called the self-perimeter of $B$.
In our approach, we replace $\omega^{(n-1)}$ in (2) by a recursive definition of $\omega^{(n-1)}_{B}(\nu)$, which is the self-volume of the unit n1n-1 dimensional ball $B\cap\Sigma^{n-1}(\nu)$ congruent with the induced norm $\|\cdot\|_{B\cap\Sigma^{n-1}(\nu)}$ on $\Sigma^{n-1}(\nu)$. See Definition 4.1 below.
@@
$P(K)=\int_{\partial K}\frac{\omega^{(n-1)}_{B}(\nu_{x})}{{\cal H}_{n-1}(B\cap\Sigma^{n-1}(\nu_{x}))}d{\cal H}_{n-1}(x)\ ,$
while the self perimeter $P(B)$ of $B\subset\mathbb{R}^{n}$ itself and its self-volume $\omega^{(n)}_{B}$ are
$P(B)=\int_{\partial B}\frac{\omega^{(n-1)}_{B}(\nu_{x})}{{\cal H}_{n-1}(B\cap\Sigma^{n-1}(\nu_{x}))}d{\cal H}_{n-1}(x)=n\omega^{(n)}_{B}\$.
This definition coincides with Busemann’s definition for $n=2$ since $\omega^{(1)}=2$ is independent of any direction $\nu\in\mathbb{R}^{2}$. However, it replaces Busemann’s preservation of the Euclidean volume of the unit ball with the preservation of the Euclidean ratio $P(B)/\omega_{B}^{(n)}=n$ of perimeter/volume of the unit ball in $n$ dimensions (Definition 4.1).
As we will show, the exact cancellation of the Jacobian in this formulation ensures that this definition of volume is strictly preserved under origin-preserving affine transformations. We extend the definition of $\omega_{B}^{(n)}$ to non-centrally symmetric convex bodies $B$, calculate explicitly the self-volume/perimeter of the $n$-simplex $\Delta^{n}$, and prove the strict convexity of $p\mapsto\omega^{(n)}_{B-p}$. We pose the following:
Conjecture 1. The function $p\in Int(B)\mapsto\omega^{(n)}_{B-p}$ is strictly convex for any any convex set $B\subset\mathbb{R}^{n}$. In particular there exists a unique minimizer $p^{*}\in int(B)$.
Conjecture 1 is valid for any convex $B\subset\mathbb{R}^{2}$ (cf. Theorem 3.4 below). In Section 4.3 we show that this conjecture is valid of any convex body in $\mathbb{R}^{n}$ composed of Cartesian products of simplices, 1D segments, and 2D convex sets. At the end of Section 4 we pose the following (see Section 4.4):
Conjecture 2.
  • The CCS (convex centrally symmetric) body in $\mathbb{R}^{n}$ of maximal self-volume is the $n$-cube $B=[-1,1]^{n}$, $\omega^{(n)}_{B}=2^{n}$.
  • For any convex set $B\subset\mathbb{R}^{n}$ containing the origin, $\omega^{(n)}_{B}\leq\frac{(n+1)^{n}}{n!}=\omega^{(n)}_{\Delta^{n}}$, where $\Delta^{n}$ is the $n$-simplex and $0\in\Delta^{n}$ is its centeroid.
@@ 3 Self-perimeter in the 2-plane
where $ds$ is the Euclidean length of the boundary of $B$, and $\vec{\tau}(s)$ is the tangent vector to the boundary at $s\in\partial B$. An explicit expression in terms of an integral on the Euclidean circle $\mathbb{S}^{1}$ is:
@@
Hence, $r_{K,p}(v)$ is a concave function of $p$.
@@
Since the self-perimeter $P(p)$ is an integral (a continuous sum) of strictly convex functions over the boundary $\partial K$ with respect to a positive length measure $ds$, the function $P(p)$ itself is strictly convex on $\text{int}(K)$.
@@ 4.1 Examples
In this example, we study the self-volume of a non-centrally symmetric ball. Let $\Delta^{n}$ be an $n$-dimensional simplex with $n+1$ vertices $v_{1},\ldots,v_{n+1}\in\mathbb{R}^{n}$. Any interior point $p\in\text{int}(\Delta^{n})$ can be uniquely represented by its strictly positive barycentric coordinates $\lambda=(\lambda_{1},\dots,\lambda_{n+1})$, where $p=\sum_{i=1}^{n+1}\lambda_{i}v_{i}$ and $\sum_{i=1}^{n+1}\lambda_{i}=1$, with $\lambda_{i}>0$ for all $i$. Let $F_{i}$ denote the (n1)(n-1)-dimensional face opposite to the $i$-th vertex.
@@
where $\{k_{1},\ldots,k_{n-1}\}\subset\{1,\ldots,n+1\}$ are n1n-1 different indices ($k_{i}\neq k_{j}$ for $i\neq j$) and $\lambda_{1},\ldots,\lambda_{n+1}$ are the barycentric coordinates of $p$.
@@ 4.2 Convexity of self-volume
The self-volume $\omega^{(n)}(\Delta^{n}-p)$ is a strictly convex function of the center $p$. In particular, there exists a unique $p^{*}\in\Delta^{n}$ for which $\omega^{(n)}(\Delta^{n}-p^{*})$ attains its minimum, and $\lambda_{i}=\frac{1}{n+1}$, $i=1,\ldots,n+1$ are the barycentric coordinates of $p^{*}$.
@@
Let $n=1$ and $m=1$. Both $A$ and $B$ are one-dimensional segments, denoted $I_{1}$ and $I_{2}$. The Cartesian product $I_{1}\times I_{2}$ is a 2D rectangle. The boundary $\partial(I_{1}\times I_{2})$ decomposes into two parts: $(\partial I_{1}\times I_{2})\cup(I_{1}\times\partial I_{2})$. For the faces $\partial I_{1}\times I_{2}$, the normal vectors are parallel to $I_{1}$. The orthogonal cross-section is $I_{2}$. The Euclidean length of these two faces is $2|I_{2}|$. The contribution to the self-perimeter is:
@@
By the recursive definition, the remaining integral is exactly the self-perimeter of $A$, denoted $P^{(n)}(A)$. Since $P^{(n)}(A)=n\omega^{(n)}_{A}$, we obtain:
@@ 5.1 Background and Motivation
In classical convex geometry, the Minkowski and Alexandrov problems seek to reconstruct a convex body from its standard Euclidean surface area measure or integral Gauss curvature, respectively. In this framework, we introduce a novel variation: an Alexandrov-type problem formulated with respect to the self-surface measure of a centrally symmetric convex set. Instead of relying on the Euclidean surface area, the perimeter and volume are defined intrinsically. For a centrally symmetric convex set $K\subset\mathbb{R}^{n}$, the self-perimeter is constructed recursively by integrating over the essential boundary, normalized by the self-volume of the (n1)(n-1)-dimensional cross-sections. The inverse problem therefore asks: Given a measure $\mu$ on $S^{n-1}$, does there exist a centrally symmetric convex set $B\subset\mathbb{R}^{n}$ whose self-surface measure is exactly $\mu$?
@@
$\int_{\{s\in\mathbb{S}^{1},\phi^{(p)}(s)+\phi_{0}>0\}}\sqrt{\phi^{(p)}(s)+\phi_{0}}\,ds-\int_{\{s\in\mathbb{S}^{1},\phi^{(p)}(s)+\phi_{0}<0\}}\sqrt{|\phi^{(p)}(s)+\phi_{0}|}\,ds=0.$ (20)
@@ 6 Conclusion
In this paper, we introduced a natural, recursive definition for volume and surface area in $n$-dimensional normed spaces which preserves the Euclidean ratio $P(B)/\omega^{(n)}_{B}(B)=n$ of perimeter to volume of the unit ball $B$. We demonstrated that this construction is invariant under origin-preserving affine transformations, satisfying fundamental geometric requirements for measures in Minkowski spaces [1, 5].
We extend this definition also for non centrally symmetric convex sets $B\subset\mathbb{R}^{n}$ and prove that the self volume of a Cartesian products of sets is the product of their self-volume. As a result we verified that the function $p\mapsto V(B-p)$ is convex for a class of convex bodies defined by a Cartesian products of simplexes, 2D convex sets and 1-D intervals. In particular we obtain the existence of an optimal center for such sets for which the self-volume is minimal.
Finally, we explore an Alexandrov-type problem for the associated self-surface measure [4]. Through a perturbative analysis around the Euclidean disc, we identified a rigidity result for the surface measure $\mu$, revealing a significant obstruction to solvability: a necessary condition for the existence of a solution is the absence of certain harmonic components in the density perturbation. These findings open new avenues for research, particularly regarding the global solvability of the inverse problem and the extremal values of the self-volume in higher dimensions.